# Adaptive Learning in Spatial Agent-Based Models for Climate Risk Assessment: A Geospatial Framework with Evolutionary Economic Agents

### v1

@@ Abstract
Climate risk assessment requires modelling complex interactions between spatially heterogeneous hazards and adaptive economic systems. We present a novel geospatial agent-based model that integrates climate hazard data with evolutionary learning for economic agents. Our framework combines Mesa-based spatial modelling with CLIMADA climate impact assessment, introducing adaptive learning behaviours that allow firms to evolve strategies for budget allocation, pricing, wages, and risk adaptation through fitness-based selection and mutation. We demonstrate the framework using riverine flood projections under RCP8.5 until 2100, showing that evolutionary adaptation enables firms to converge with baseline (no hazard) production levels after decades of disruption due to climate stress. Our results reveal systemic risks where even agents that are not directly exposed to floods face impacts through supply chain disruptions, with the end-of-century average price of goods 5.6% higher under RCP8.5 compared to the baseline. This open-source framework provides financial institutions and companies with tools to quantify both direct and cascading climate risks while evaluating cost-effective adaptation strategies.
@@ 2.3 Evolutionary learning system
Performance Memory: Each firm maintains a 10-step rolling window tracking money, production, capital stock, and limiting factors (labor/capital/input constraints).
Fitness Function: Combines four components with fixed weights: 1) Growth rate (40%): Money growth rate with diminishing returns via tanh; 2) Production stability (30%): One minus the coefficient of variation in production (rewards consistent production with lower variability); 3) Survival bonus (20%): Longevity reward that increases linearly up to 20 steps; 4) Resource balance (10%): Diversity of limiting factors across labor, capital, and inputs.
Mutation Process: Strategy parameters mutated every 5 steps with 30% probability per parameter, using Gaussian noise with standard deviation of 2.5% (if fitness improved), 5% (initial), or 10% (if fitness declined), implementing simple hill-climbing.
Population Dynamics: Failed firms are replaced by offspring of successful firms, with up to 25% population replacement per step after an initial 5-step establishment period. Failed firms are defined as firms that have less money than the minimum survival amount (set to 1 monetary unit), or those that are in consistent decline such that their wealth is reduced by more than 50% over 5 time-steps.
@@ 3 Results and Discussion
These results highlight the importance of geospatial agent-based modelling in capturing supply chain disruptions due to climate risk. Agents throughout the network are exposed to risks even if they are not directly affected by acute hazards. While hazards affect the capital, inventory, and productivity of firms, households face several system risks, including higher unemployment rates and higher inflation. Firms also experience indirect risks, as evident by the near doubling of the percentage of input-limited firms under RCP8.5 compared to baseline in 2050 (Figure 1h vs. 1i). However, firms are able to lower these risks by adapting to evolving climate stresses, as shown by the convergence of firm production and wealth under RCP8.5 with those of the baseline by the end of the century.
Figure 1: Agent trajectories under baseline (no hazard) and RCP8.5 riverine flooding scenarios. Com-B and Man-B refer to commodity and manufacturing firms. B and H labels refer to Baseline and Hazard (RP8.5) scenarios, respectively.
@@ 4 Conclusions
This study was funded by Pluripotent Technologies Inc.

### v2

@@ Abstract
Climate risk assessment requires modelling complex interactions between spatially heterogeneous hazards and adaptive economic systems. We present a novel geospatial agent-based model that integrates climate hazard data with evolutionary learning for economic agents. Our framework combines Mesa-based spatial modelling with CLIMADA climate impact assessment, introducing adaptive learning behaviours that allow firms to evolve strategies for budget allocation, pricing, wages, and risk adaptation through fitness-based selection and mutation. We demonstrate the framework using riverine flood projections under RCP8.5 until 2100, showing that evolutionary adaptation enables firms to converge with baseline (no hazard) production levels after decades of disruption due to climate stress. Our results reveal systemic risks where even agents that are not directly exposed to floods face impacts through supply chain disruptions, with the end-of-century average price of goods 5.6% higher under RCP8.5 compared to the baseline in our illustrative economic network. This open-source framework provides financial institutions and companies with tools to quantify both direct and cascading climate risks while evaluating cost-effective adaptation strategies.
@@ 2.3 Evolutionary learning system
Performance Memory: Each firm maintains a 10-step rolling window tracking money, production, capital stock, and limiting factors (labor/capital/input constraints), with fitness evaluation based on the most recent 5 steps.
Fitness Function: Combines four components with fixed weights: 1) Growth rate (40%): Money growth rate with diminishing returns via tanh transformation; 2) Production stability (30%): One minus the coefficient of variation in production (rewards consistent production with lower variability); 3) Survival bonus (20%): Longevity reward that increases linearly up to 20 steps; 4) Resource balance (10%): Diversity of limiting factors across labor, capital, and inputs.
Individual Adaptation: Each living firm independently evaluates and mutates its strategy parameters every 5 steps. Mutation strength adapts based on fitness changes: parameters have a 30% probability of mutation with Gaussian noise at standard deviation of 2.5% of current value (if fitness improved), 5% (baseline), or 10% (if fitness declined), implementing adaptive hill-climbing.
Evolutionary Replacement: Up to a quarter of failed firms are replaced every 10 steps – defined as firms that have less money than the minimum survival amount or their wealth reduced by more than 50% over 5 time-steps. Offspring are created from fitness-weighted selection of successful firms, inheriting parent strategy with a 50% probability of mutation per parameter with Gaussian noise at 10% standard deviation.
@@ 3 Results and Discussion
These results highlight the importance of geospatial agent-based modelling in capturing systemic disruptions due to climate risk. Agents throughout the network are exposed to risks even if they are not directly affected by acute hazards. While hazards affect the capital, inventory, and productivity of firms, households face several system risks, including higher unemployment rates and higher inflation. Firms also experience indirect risks, as evident by the near doubling of the percentage of input-limited firms under RCP8.5 compared to baseline in 2050 (Figure 1h vs. 1i). However, firms are able to lower these risks by adapting to evolving climate stresses, as shown by the convergence of firm production and wealth under RCP8.5 with those of the baseline by the end of the century.
Figure 1: Agent trajectories under Baseline (no hazard) and Hazard (RCP8.5 riverine flooding) scenarios.
@@ 4 Conclusions
This work was supported by Quantile Labs.